What is the pH of a 0.10M solution of methylamine (:NH2CH3) From Table 15.8, Kb = 4.4 x 10-4.

Please show me how to work it out, thank you!
Answer #1

ans)

Here

CH3NH2 + H2O -----> CH3NH3+ + OH-

0.10 ............................x.................x

Kb = [Ch3NH3+][OH-]/[CH3NH2]

4.4*10^-4 = (x)*(x)/0.10

x = [OH-] = 0.663*10^-2

As we know that

pH = 14 - p[OH]

= 14 - (- log[OH-])

= 14 + log[OH-]

= 14 + log(0.663*10^-2)

= 11.82